OpenStudy (konradzuse):

Find the average value of f(x) = 48-12x^2 of the interval -5<=x<=5

OpenStudy (konradzuse):

This sounds like an integration question, but I don't think I've seen integration start iwht y or f(x) so I was thinking derivation? Just confused how to start this.

OpenStudy (goformit100):

@ParthKohli

OpenStudy (anonymous):

average value is \[\frac{1}{10}\int_{-5}^5 (48-12x^2)dx\]

OpenStudy (konradzuse):

how did you figure that out?

OpenStudy (anonymous):

that is the definition of average value integral divided by the length of the path but i am going to guess that it is zero in any case since this is a parabola with axis of symmetry as the \(y\) axis

OpenStudy (anonymous):

scratch that remark it is wrong

OpenStudy (konradzuse):

http://archives.math.utk.edu/visual.calculus/5/average.1/index.html From this it makes sense 1/(5--5)

OpenStudy (anonymous):

it would be zero if the function was odd, sorry ignore that last remark

OpenStudy (konradzuse):

actually no -5-5 = 0

OpenStudy (konradzuse):

1/0 is undefined though?

OpenStudy (anonymous):

it is \(\frac{1}{b-a}\) in your case \(\frac{1}{5-(-5)}=\frac{1}{10}\) but you should think of it simply as the length of the path

OpenStudy (anonymous):

from minus 5 to 5 is ten

OpenStudy (konradzuse):

oh wait b is 5 lol I'm confusing myself.

OpenStudy (anonymous):

that is why it is called "average value" it is analogous to the mean

OpenStudy (anonymous):

if you have a bunch of numbers to find the average you add up and divide by how many you have this is the continuous version of the same thing. adding, then take the limit is the integral, how many you have is the length of the interval so it really is an average but with continuous data instead of discrete

OpenStudy (konradzuse):

yeah I would have thought for 2 values it would have been 2/10 but this formula is different. Do I now do the integral, or am I done?